# Particle-Grid Neural Dynamics for Learning Deformable Object Models from RGB-D Videos

### v1

@@ III-E Planning
We apply the Model-Predictive Path Integral (MPPI) [50] trajectory optimization algorithm to minimize the cost and to synthesize the robot’s actions. During deployment, we perform online, iterative planning to achieve closed-loop control.
@@ IV Experiments
MPM [12] $\mathcal{J}$-Score / IoU $\uparrow$ $40.5_{\pm 20.3}$ $12.5_{\pm 9.5}$ $37.5_{\pm 14.8}$- $34.3_{\pm 5.5}$ $42.9_{\pm 10.3}$
GBND [58] $56.7_{\pm 13.9}$ $19.2_{\pm 15.6}$ $42.7_{\pm 10.5}$ $\mathbf{83.0}_{\mathbf{\pm 5.5}}$ $\mathbf{78.0}_{\mathbf{\pm 82.8}}$ $49.8_{\pm 13.9}$
@@
MPM [12] $\mathcal{F}$-Score $\uparrow$ $28.0_{\pm 11.3}$ $37.0_{\pm 14.2}$ $40.4_{\pm 12.2}$- $13.1_{\pm 5.9}$ $54.3_{\pm 12.1}$
GBND [58] $32.2_{\pm 18.2}$ $41.9_{\pm 18.9}$ $35.8_{\pm 11.2}$ $\mathbf{74.0}_{\mathbf{\pm 8.8}}$ $54.6_{\pm 13.4}$ $60.1_{\pm 16.1}$
@@
MPM [12] LPIPS $\downarrow$ $0.141_{\pm 0.060}$ $0.052_{\pm 0.018}$ $0.103_{\pm 0.085}$- $0.145_{\pm 0.020}$ $0.059_{\pm 0.020}$
GBND [58] $0.120_{\pm 0.055}$ $0.042_{\pm 0.018}$ $0.081_{\pm 0.023}$ $0.106_{\pm 0.039}$ $0.097_{\pm 0.019}$ $0.052_{\pm 0.015}$
Particle $0.109_{\pm 0.048}$ $0.044_{\pm 0.032}$ $0.072_{\pm 0.024}$ $0.082_{\pm 0.031}$ $0.069_{\pm 0.019}$ $0.054_{\pm 0.020}$
@@
TABLE II: Quantitative Results on 3D Action-Conditioned Video Prediction. We compared our method on 3D action-conditioned video prediction quality with MPM [12], GBND [58], and particle-based baselines. The $\mathcal{J}$-Score/IoU and the $\mathcal{F}$-Score measures mask similarities and the LPIPS score measures appearance-wise similarities between predicted frames and ground truth video recordings. We report the mean and standard deviation of the prediction error over a 3-second horizon. The best results are highlighted in bold and blue.
@@ IV-A Experiment Setup
  • MPM-based deformable object simulation [12, 29]: This baseline assumes a hyperelastic material with an unknown uniform Young’s modulus and friction coefficient with the tabletop. Parameter identification is performed via gradient descent.
  • Graph-Based Neural Dynamics (GBND) [58]: This model represents objects using subsampled sparse vertices, along with the robot end-effectors, and employs a Graph Neural Network (GNN) to predict particle motions.
@@ IV-B Dynamics Learning and Prediction
[…]
@@ IV-D Category-Level Model
The results in Fig. 5 show that our model achieves lower prediction errors than the GBND baseline across both categories and for both seen and unseen instances. Notably, for the cloth category, the baseline method exhibits a significantly performance drop on unseen instances, whereas our method keeps a relatively low error, demonstrating better generalization to novel objects at test time.
@@ IV-E Action-Conditioned Video Prediction
For action-conditioned video prediction, we use the predicted point cloud trajectories to interpolate Gaussian kernel transformations using LBS [45]. We reconstruct Gaussians from 4 input views using Gaussian Splatting [16]. The Gaussians are trained with a segmentation mask loss, following previous works [58, 28]. Videos are rendered with a fixed input camera pose. The video prediction quality is assessed using mask-based metrics, including $\mathcal{J}$-Score (IoU), $\mathcal{F}$-Score (contour matching accuracy), and the image-based metric LPIPS.
The resulting metrics are shown in Table II. Our approach achieves the best overall performance. For categories with relatively large objects, for instance boxes and paper bags, we observe that spiky Gaussian reconstructions often negatively impact mask prediction performance, especially when objects undergo significant deformation. The higher mask alignment scores in GBND and Particle baselines are largely due to inadequate particle motion predictions.
@@ V Limitations
While we have demonstrated that Particle-Grid Neural Dynamics can model diverse types of deformable objects, the current framework has several limitations: (i) The current formulation assumes a fixed number of particles during iterative rollout, making it inapplicable to scenarios involving the appearance or disappearance of particles. This limitation could be addressed by correcting the particle sets with new observations or modeling the per-frame visibility of particles. (ii) The model implicitly infers an object’s physical properties from its point cloud and short-term motion history. While this is sufficient for modeling a single object instance or multiple instances with distinct shapes and physical properties, a more systematic approach to modeling physical properties is needed for interpretable identification and adaptation at test time. This could involve learning a parameter-conditioned neural dynamics model [57]. (iii) Training the model and applying it to video prediction depend on accurate predictions from computer vision models such as Segment-Anything [18], CoTracker [15], and Gaussian Splatting [16]. Failures in these perception and reconstruction models could negatively impact our method’s performance.
@@ Acknowledgment
Figure 10: Experiment Setup. (a) Our robot workspace includes four calibrated RGB-D cameras positioned at each corner of the table, along with a GELLO [51] system for teleoperating the dual xArm 7 robotic arms equipped with parallel grippers. (b) Our experiments involved six types of materials: (i) a paper bag, (ii) a stuffed animal, (iii) three varieties of bread, (iv) three boxes of different shapes, (v) eight cloth pieces varying in fabric type and size, and (vi) six ropes differing in length, thickness, and stiffness.
[…]
1Contents
@@ A.1 Data Collection
or each camera view, we apply Grounded-SAM-2 [38, 18, 39] to extract object masks. Using text prompts containing object descriptions, Grounded-SAM-2 detects and segments object in the first frame of a sequence and then propagates the mask across subsequent frames. To apply the model, we partition the recording into non-overlapping clips, each containing 720 frames (i.e., 24 seconds).
@@
We use CoTracker [15] as the tracking model due to its stable and robust performance in 2D tracking. Each camera view’s recording is parsed into clips of 30 frames (i.e., 1 second) with 15 overlapping frames. Within each clip, we crop objects from the video based on the segmentation mask and fill the background with black pixels. Using the cropped the clip as input, CoTracker predicts a set of 2D trajectories, initialized from uniformly sampled grid locations in the first frame of the clip. We then interpolate each pixel’s 3D velocity using the predicted trajectories and the corresponding depth image, as described by the following equation:
@@ A.2 Video Prediction
Equivalently, $\mathbf{r}_{p}$ represents the optimal rotation that transforms the nearest particles of $p$ to their new positions. Using the estimated rotations, we perform Linear Blend Skinning (LBS) [45] to interpolate the transformations. For a Gaussian kernel $i$ with $\mu_{i,t}\in\mathbf{X}_{\text{GS},t}$ and $q_{i,t}\in\mathbf{R}_{\text{GS},t}$, we denote its $k_{\text{LBS}}$ nearest particles to be $\mathcal{N}(i)$. We calculate the interpolated positions and quaternions, $\hat{\mu}_{i,t+\Delta t}$ and $\hat{q}_{i,t+\Delta t}$, using the following equations:
@@ A.3 Model-Based Planning
For the four manipulation experiments, the planning settings are detailed as follows. The cloth lifting and box closing tasks are bimanual, while the rope manipulation and plush toy relocation tasks involve only a single arm. For the latter two tasks, we do not consider gripper rotations, so the action space consists of the 3-DoF $(x,y,z)$ position of the end effector in the task space. For the two bimanual tasks, we account for 3D rotations, resulting in an action space that includes the translation and rotation of both grippers, totaling 12 degrees of freedom. In each iteration of the MPPI [50] algorithm, we sample $N$ delta actions from a normal distribution:
@@
[…]
@@ B.1 Experiment Setup
Our robot setup and experimental objects are shown in Fig. 10. Our experiments, including real-world data collection and manipulation tasks, are conducted in a workspace (Fig. 10a) equipped with four calibrated RealSense D455 cameras. These cameras are positioned at the corners of the table to capture RGB-D images at 30Hz with a resolution of 848x480. Additionally, we have integrated the GELLO [51] teleoperation system with dual UFACTORY xArm 7 robotic arms, each with 7 degrees of freedom and xArm’s parallel grippers, enabling a human operator to collect bi-manual random interaction data.
@@ B.21 MPM
We follow previous works [53, 29] for the implementation and parameter optimization of MPM. In the MPM algorithm, an object is discretized into Lagrangian particles that carry dynamic information such as position, velocity, strain, and stress. The physical properties of the object are described by the material’s elasticity and plasticity models, known as constitutive models. Different object types require different constitutive models, and objects within the same type may have varying parameters. We assume that the materials considered in our work are hyperelastic, and thus we use the fixed corotated elasticity model with an identity plasticity model. The optimizable parameters in the models include Young’s modulus $E$, which characterizes the stiffness or stress in response to deformations, and the Poisson’s ratio $\nu$, which is kept fixed. We also optimize the friction coefficient $\mu$ between the object and surface. Since we assume uniform material properties across the objects, the optimizable parameters $E$ and $\mu$ are shared across all particles.
@@ B.22 GBND
The GBND method defines the environment state as a graph: $z_{t}\overset{\underset{\mathrm{\Delta}}{}}{=}\mathcal{G}_{t}=(\mathcal{V}_{t},\mathcal{E}_{t})$, where $\mathcal{V}_{t}$ is the vertex set representing object particles, and $\mathcal{E}_{t}$ denotes the edge set representing interacting particle pairs at time step $t$. Given dense particles of object and robot end-effector pose, we apply farthest point sampling [32] to downsample object into sparse vertices, while representing the robot gripper with a single particle. For a vertex $v_{i,t}\in\mathcal{V}_{t}$, we incorporate the history positions $\mathbf{x}_{i,t-h\Delta t:t}$ to implicitly capture velocity information, along with vertex attributes $\mathbf{c}_{i,t}$ to indicate whether the particle corresponds to the object or the robot end-effector. The final input is the concatenated feature vector $(\mathbf{x}_{i,t-h\Delta t:t},\mathbf{c}_{i,t})$. In practice, we set $h=2$ for all materials. For end-effector particles, the action $\mathbf{u}_{t}$, represented as delta gripper transformation, is also included in the input feature. The edges between particles are dynamically constructed over time by identifying the nearest neighbors of each particle within a predefined radius. Additionally, we impose a limit on the maximum number of edges that any single node can have, which we empirically set to 5.
After constructing the graph, we instantiate the dynamics model as graph neural networks (GNNs) that predict the evolution of the graph representation $z_{t}$. The graph design follows previous works [19, 58]. To control the accumulation of dynamics prediction errors, we supervise the model’s predictions over $K=4$ steps. Formally, we train the model using following the loss function:
@@
where $\hat{\mathbf{x}}_{i,t}$ and $\mathbf{x}_{i,t}$ represent the predicted and ground truth positions of the $i^{\text{th}}$ vertex, respectively. The first term in the loss is the mean squared error, while the second term is an edge length regularization term, which has been shown to be effective for objects with complex shapes [58]. Empirically, we set $\lambda=0.1$.
@@ B.3 Method Implementation
Following previous works [53, 29], we implement the grid velocity editing and G2P processes with NVIDIA Warp [30], which accelerates the differentiable process on GPU. Training our model typically takes around 8 hours on a single NVIDIA 4090 GPU, using a batch size of 32.
[…]

### v2

@@ III-E Planning
We apply the Model-Predictive Path Integral (MPPI) [williams2017model] trajectory optimization algorithm to minimize the cost and to synthesize the robot’s actions. During deployment, we perform online, iterative planning to achieve closed-loop control.
@@ IV Experiments
MPM [hu2018moving] $\mathcal{J}$-Score / IoU $\uparrow$ $40.5_{\pm 20.3}$ $12.5_{\pm 9.5}$ $37.5_{\pm 14.8}$ - $34.3_{\pm 5.5}$ $42.9_{\pm 10.3}$
GBND [zhang2024dynamic] $56.7_{\pm 13.9}$ $19.2_{\pm 15.6}$ $42.7_{\pm 10.5}$ $\mathbf{83.0}_{\mathbf{\pm 5.5}}$ $\mathbf{78.0}_{\mathbf{\pm 82.8}}$ $49.8_{\pm 13.9}$
@@
MPM [hu2018moving] $\mathcal{F}$-Score $\uparrow$ $28.0_{\pm 11.3}$ $37.0_{\pm 14.2}$ $40.4_{\pm 12.2}$ - $13.1_{\pm 5.9}$ $54.3_{\pm 12.1}$
GBND [zhang2024dynamic] $32.2_{\pm 18.2}$ $41.9_{\pm 18.9}$ $35.8_{\pm 11.2}$ $\mathbf{74.0}_{\mathbf{\pm 8.8}}$ $54.6_{\pm 13.4}$ $60.1_{\pm 16.1}$
@@
MPM [hu2018moving] LPIPS $\downarrow$ $141_{\pm 0.060} 0.0$$052_{\pm 0.018}$ $0.103_{\pm 0.085}$ - $145_{\pm 0.020} 0.0$$059_{\pm 0.020}$
GBND [zhang2024dynamic] $120_{\pm 0.055} 0.0$$42_{\pm 0.018} 0.0$$081_{\pm 0.023}$ $106_{\pm 0.039} 0.0$$97_{\pm 0.019} 0.0$$052_{\pm 0.015}$
Particle $109_{\pm 0.048} 0.0$$44_{\pm 0.032} 0.0$$72_{\pm 0.024} 0.0$$82_{\pm 0.031} 0.0$$69_{\pm 0.019} 0.0$$054_{\pm 0.020}$
@@
TABLE II: Quantitative Results on 3D Action-Conditioned Video Prediction. We compared our method on 3D action-conditioned video prediction quality with MPM [hu2018moving], GBND [zhang2024dynamic], and particle-based baselines. The $\mathcal{J}$-Score/IoU and the $\mathcal{F}$-Score measures mask similarities and the LPIPS score measures appearance-wise similarities between predicted frames and ground truth video recordings. We report the mean and standard deviation of the prediction error over a 3-second horizon. The best results are highlighted in bold and blue.
@@ IV-A Experiment Setup
  • MPM-based deformable object simulation [hu2018moving, ma2023learning]: This baseline assumes a hyperelastic material with an unknown uniform Young’s modulus and friction coefficient with the tabletop. Parameter identification is performed via gradient descent.
  • Graph-Based Neural Dynamics (GBND) [zhang2024dynamic]: This model represents objects using subsampled sparse vertices, along with the robot end-effectors, and employs a Graph Neural Network (GNN) to predict particle motions.
@@ IV-B Dynamics Learning and Prediction
We evaluate accuracy using a held-out set of robot-object interactions. The interaction videos are divided into 3-second clips, and the dynamics rollout accuracy is assessed using 3D point cloud metrics that compute the distance between predicted particle positions and ground truth future points. The metrics include Mean Distance Error (MDE, the average distance between the corresponding predicted and ground truth points), Chamfer Distance (CD, the average distance between pairs of nearest neighbors between two point clouds), and Earth Mover’s Distance (EMD, the average distance between two point clouds according to an optimal correspondence). All metrics are calculated with  $\text{\,}\mathrm{m}$ as the unit of measurement. For the box category, we use the Robot Particles control method representation since it is not directly compatible with MPM. Therefore, we omit MPM from the box comparison.
@@ IV-D Category-Level Model
The results in Fig. 5 show that our model achieves lower prediction errors than the GBND baseline across both categories and for both seen and unseen instances. Notably, for the cloth category, the baseline method exhibits a significant performance drop on unseen instances, whereas our method keeps a relatively low error, demonstrating better generalization to novel objects at test time.
@@ IV-E Action-Conditioned Video Prediction
For action-conditioned video prediction, we use the predicted point cloud trajectories to interpolate Gaussian kernel transformations using LBS [10.1145/1276377.1276478]. We reconstruct Gaussians from 4 input views using Gaussian Splatting [kerbl3Dgaussians]. The Gaussians are trained with a segmentation mask loss, following previous works [zhang2024dynamic, luiten2023dynamic]. Videos are rendered with a fixed input camera pose. The video prediction quality is assessed using mask-based metrics, including $\mathcal{J}$-Score (IoU), $\mathcal{F}$-Score (contour matching accuracy), and the image-based metric LPIPS.
The resulting metrics are shown in Table II. Our approach achieves the best overall performance. For categories with relatively large objects, for instance, boxes and paper bags, we observe that spiky Gaussian reconstructions often negatively impact mask prediction performance, especially when objects undergo significant deformation. The higher mask alignment scores in GBND and Particle baselines are largely due to inadequate particle motion predictions.
@@ V Limitations
While we have demonstrated that Particle-Grid Neural Dynamics can model diverse types of deformable objects, the current framework has several limitations: (i) The current formulation assumes a fixed number of particles during iterative rollout, making it inapplicable to scenarios involving the appearance or disappearance of particles. This limitation could be addressed by correcting the particle sets with new observations or modeling the per-frame visibility of particles. (ii) The model implicitly infers an object’s physical properties from its point cloud and short-term motion history. While this is sufficient for modeling a single object instance or multiple instances with distinct shapes and physical properties, a more systematic approach to modeling physical properties is needed for interpretable identification and adaptation at test time. This could involve learning a parameter-conditioned neural dynamics model [zhang2024adaptigraph]. (iii) Training the model and applying it to video prediction depend on accurate predictions from computer vision models such as Segment-Anything [kirillov2023segment], CoTracker [karaev24cotracker3], and Gaussian Splatting [kerbl3Dgaussians]. Failures in these perception and reconstruction models could negatively impact our method’s performance.
@@ Acknowledgment
Figure 10: Experiment Setup. (a) Our robot workspace includes four calibrated RGB-D cameras positioned at each corner of the table, along with a GELLO [wu2023gello] system for teleoperating the dual xArm 7 robotic arms equipped with parallel grippers. (b) Our experiments involved six types of materials: (i) a paper bag, (ii) a stuffed animal, (iii) three varieties of bread, (iv) three boxes of different shapes, (v) eight cloth pieces varying in fabric type and size, and (vi) six ropes differing in length, thickness, and stiffness.
Contents
@@ A.1 Data Collection
or each camera view, we apply Grounded-SAM-2 [ravi2024sam2segmentimages, kirillov2023segment, ren2024grounded] to extract object masks. Using text prompts containing object descriptions, Grounded-SAM-2 detects and segments object in the first frame of a sequence and then propagates the mask across subsequent frames. To apply the model, we partition the recording into non-overlapping clips, each containing 720 frames (i.e., 24 seconds).
@@
We use CoTracker [karaev24cotracker3] as the tracking model due to its stable and robust performance in 2D tracking. Each camera view’s recording is parsed into clips of 30 frames (i.e., 1 second) with 15 overlapping frames. Within each clip, we crop objects from the video based on the segmentation mask and fill the background with black pixels. Using the cropped the clip as input, CoTracker predicts a set of 2D trajectories, initialized from uniformly sampled grid locations in the first frame of the clip. We then interpolate each pixel’s 3D velocity using the predicted trajectories and the corresponding depth image, as described by the following equation:
@@ A.2 Video Prediction
Equivalently, $\mathbf{r}_{p}$ represents the optimal rotation that transforms the nearest particles of $p$ to their new positions. Using the estimated rotations, we perform Linear Blend Skinning (LBS) [10.1145/1276377.1276478] to interpolate the transformations. For a Gaussian kernel $i$ with $\mu_{i,t}\in\mathbf{X}_{\text{GS},t}$ and $q_{i,t}\in\mathbf{R}_{\text{GS},t}$, we denote its $k_{\text{LBS}}$ nearest particles to be $\mathcal{N}(i)$. We calculate the interpolated positions and quaternions, $\hat{\mu}_{i,t+\Delta t}$ and $\hat{q}_{i,t+\Delta t}$, using the following equations:
@@ A.3 Model-Based Planning
For the four manipulation experiments, the planning settings are detailed as follows. The cloth lifting and box closing tasks are bimanual, while the rope manipulation and plush toy relocation tasks involve only a single arm. For the latter two tasks, we do not consider gripper rotations, so the action space consists of the 3-DoF $(x,y,z)$ position of the end effector in the task space. For the two bimanual tasks, we account for 3D rotations, resulting in an action space that includes the translation and rotation of both grippers, totaling 12 degrees of freedom. In each iteration of the MPPI [williams2017model] algorithm, we sample $N$ delta actions from a normal distribution:
@@
We then evaluate the sampled action trajectories using the learned dynamics model and calculate the cost based on Chamfer Distance (Eq. 16). For the cloth lifting task, to prevent damage to the cloth, we include an additional penalty term with an indicator function that penalizes trajectories where the distance between the grippers exceeds $L=$0.6\text{\,}\mathrm{m}$\$.
@@ B.1 Experiment Setup
Our robot setup and experimental objects are shown in Fig. 10. Our experiments, including real-world data collection and manipulation tasks, are conducted in a workspace (Fig. 10a) equipped with four calibrated RealSense D455 cameras. These cameras are positioned at the corners of the table to capture RGB-D images at 30Hz with a resolution of 848x480. Additionally, we have integrated the GELLO [wu2023gello] teleoperation system with dual UFACTORY xArm 7 robotic arms, each with 7 degrees of freedom and xArm’s parallel grippers, enabling a human operator to collect bi-manual random interaction data.
@@ B.21 MPM
We follow previous works [xie2023physgaussian, ma2023learning] for the implementation and parameter optimization of MPM. In the MPM algorithm, an object is discretized into Lagrangian particles that carry dynamic information such as position, velocity, strain, and stress. The physical properties of the object are described by the material’s elasticity and plasticity models, known as constitutive models. Different object types require different constitutive models, and objects within the same type may have varying parameters. We assume that the materials considered in our work are hyperelastic, and thus we use the fixed corotated elasticity model with an identity plasticity model. The optimizable parameters in the models include Young’s modulus $E$, which characterizes the stiffness or stress in response to deformations, and the Poisson’s ratio $\nu$, which is kept fixed. We also optimize the friction coefficient $\mu$ between the object and surface. Since we assume uniform material properties across the objects, the optimizable parameters $E$ and $\mu$ are shared across all particles.
@@ B.22 GBND
The GBND method defines the environment state as a graph: $z_{t}\overset{\underset{\mathrm{\Delta}}{}}{=}\mathcal{G}_{t}=(\mathcal{V}_{t},\mathcal{E}_{t})$, where $\mathcal{V}_{t}$ is the vertex set representing object particles, and $\mathcal{E}_{t}$ denotes the edge set representing interacting particle pairs at time step $t$. Given dense particles of object and robot end-effector pose, we apply farthest point sampling [FarthestPointSampling] to downsample object into sparse vertices, while representing the robot gripper with a single particle. For a vertex $v_{i,t}\in\mathcal{V}_{t}$, we incorporate the history positions $\mathbf{x}_{i,t-h\Delta t:t}$ to implicitly capture velocity information, along with vertex attributes $\mathbf{c}_{i,t}$ to indicate whether the particle corresponds to the object or the robot end-effector. The final input is the concatenated feature vector $(\mathbf{x}_{i,t-h\Delta t:t},\mathbf{c}_{i,t})$. In practice, we set $h=2$ for all materials. For end-effector particles, the action $\mathbf{u}_{t}$, represented as delta gripper transformation, is also included in the input feature. The edges between particles are dynamically constructed over time by identifying the nearest neighbors of each particle within a predefined radius. Additionally, we impose a limit on the maximum number of edges that any single node can have, which we empirically set to 5.
After constructing the graph, we instantiate the dynamics model as graph neural networks (GNNs) that predict the evolution of the graph representation $z_{t}$. The graph design follows previous works [li2018learning, zhang2024dynamic]. To control the accumulation of dynamics prediction errors, we supervise the model’s predictions over $K=4$ steps. Formally, we train the model using following the loss function:
@@
where $\hat{\mathbf{x}}_{i,t}$ and $\mathbf{x}_{i,t}$ represent the predicted and ground truth positions of the $i^{\text{th}}$ vertex, respectively. The first term in the loss is the mean squared error, while the second term is an edge length regularization term, which has been shown to be effective for objects with complex shapes [zhang2024dynamic]. Empirically, we set $\lambda=0.1$.
@@ B.3 Method Implementation
Following previous works [xie2023physgaussian, ma2023learning], we implement the grid velocity editing and G2P processes with NVIDIA Warp [warp2022], which accelerates the differentiable process on GPU. Training our model typically takes around 8 hours on a single NVIDIA 4090 GPU, using a batch size of 32.
In all experiments, we use a time step of $\Delta t=$0.1\text{\,}\mathrm{s}$\$. To effectively model smaller objects, such as bread, we introduce a scaling factor $s$ that scales the object before being input into the model. Specifically, we use $s=3.0$ and grid size $\delta=$1\text{\,}\mathrm{cm}$\$ for bread, $s=0.8,\delta=$2\text{\,}\mathrm{cm}$\$ for cloth, and $s=1.0,\delta=$2\text{\,}\mathrm{cm}$\$ for all other object categories. In all experiments, the particle velocities are set to zero at the start of an episode, assuming that the object starts from a static state particle. Otherwise, the velocities are computed from the difference between current and previous observations.